# Leveraging Large Language Models for Active Merchant Non-player Characters

### v2

@@ IV-A Two possible approaches
First, we employed the ZSP method in which an LLM generates negotiation dialogues without being exposed to examples. We initially considered a naïve negotiation method without using any specific tactics. This approach relied on using the pretrained knowledge of LLMs, which reflects a general understanding of the world; consequently, the generated negotiation strayed significantly from the intended game setting. Thus, we inputted 10 negotiation tactics within the prompt, as shown in Table II, which are inspired by [6, 7, 8, 9]. We evaluated three Llama variants as in the appraiser.
TABLE II: Tactics Used in the Input Prompt to Help in Negotiations
@@
Second, we used the KD method to improve the performance of smaller LLMs. By transferring the knowledge of a larger model to a smaller model, the distillation reduces computational cost while maintaining performance. We trained smaller student models on dialogue dataset generated by larger teacher model. Here, to distill the persuasiveness of teacher model, we asked them to choose appropriate tactic before generating explicit utterances. To reduce the computational demands of LLMs during training, we adopted quantization and low-rank adaptation (LoRA) [10]. Using these methods, we distilled knowledge about negotiation from Llama 3.1 405b to two smaller LLMs: Llama 3.1 8b and 3.2 1b.
IV-B Dataset
@@
We prepared a dataset by generating a negotiation dialogue for each item in the appraiser dataset. To simulate a negotiation dialogue, we used two agents: merchant and player. As the merchant, we used Llama 3.1 405B, the teacher model. As the player, we used GPT-4o with input prompts about tactics. We used different models for these two agents to (1) avoid adopting a similar reasoning process and (2) make the player as similar as possible to human beings, as reported in [11]. To simulate the diversity of real-world customers, we used a temperature value of 1.0 for GPT-4o.
@@
The negotiation procedure was as follows. Before starting a negotiation, both agents received an item description of the negotiation subject as shared information. The agents started with different prices for each item: the player wanted to buy the item at a discount of 10% to 25%. The player started the negotiation with an utterance “Hello. I’m looking for [the item name].” While an actual negotiation might contain greetings or small talk in the real world, we avoided generating messages that are less related to the negotiation to make the model focus on learning negotiations.
@@
After starting the utterance, the player and merchant negotiated through text, proceeding until the player decided whether to buy the item or not. Until the decision was made, the merchant kept persuading the player to buy the item to simulate a real-world merchant. The player could terminate the conversation by mentioning “conversation over.” To avoid long dialogues, we set the maximum number of turns to 15. As a result, we generated 2,943 conversations: 2,616 conversations from the training set items and 327 conversations from the validation set items.
IV-C Implementation detail
The negotiator was implemented in the same way as the appraiser module, with differences only in the training details. For Python libraries, we used transformers 4.46.2 [4], trl 0.12 [12], and peft 0.13 [13]. For training, we employed 4-bit quantization and gradient accumulation across four steps on a single GPU in order to reduce memory consumption.
IV-D Evaluation metrics
@@
We compared three ZSP and two KD models using three measures: persuasiveness, dominance, and agreement. First, we measured the persuasiveness of utterances generated by the negotiator. Similar to commercial behavior in the real world, the merchant NPC should effectively persuade players to buy or sell an item at a good price. Various tactics are used in an effective negotiation; so, we aimed to evaluate the effectiveness of the negotiator in using 10 different tactics to create persuasive statements. To measure persuasiveness, we employed G-Eval [14], a widely-used evaluation method using LLMs. The method directly asks GPT-4 to evaluate an input text according to given criteria. Specifically, we used a 5-point scale and averaged 20 runs following G-Eval.
@@
Second, we measured the dominance of the merchant over the player during the negotiation. This metric indicates whether the merchant holds more power in the relationship with the player, highlighting the concept of power dynamics [15]. A human merchant usually has the initiative in price negotiation, rather than giving the initiative to the customer. So, as a negotiated price increases, the merchant profits more while the player incurs greater losses. In other words, the negotiation is a type of zero-sum game. Considering such power dynamics, we measured dominance using Equation 4, where $y$, $y^{m}$, and $y^{p}$ indicate the agreed price, retail price, and price desired by the player, respectively. The equation quantifies the ratio between the profit gained by the merchant ($y_{i}-y_{i}^{p}$) and the price gap between the two agents ($y_{i}^{m}-y_{i}^{p}$).
@@
Third, we defined the agreement rate. In the real world, two agents in negotiation may not reach an agreement. We determined how frequently the models concluded negotiations with an agreement. Based on the labeled result for dominance, we estimated the number of dialogues ending with an agreement. Then, the agreement rate was computed as the ratio of the agreed dialogues to the total number of dialogues.
TABLE III: Results of Assessing Two Methods of Negotiator
@@
* Numbers in parenthesis indicate the number of utterances.
IV-E Results
Table III shows the results of our experiments, comparing ZSP and KD methods. The results show that the KD-8b performed the best and successfully used persuasive tactics. The model achieved a score of 3.99 in terms of persuasiveness, followed by ZSP-405b (3.92), ZSP-8b (3.74), KD-1b (3.65), and ZSP-1b (2.95). The ZSP-405b had the strongest dominance (0.47), followed by ZSP-8b (0.42), and KD-8b (0.40). The ZSP-8b had the highest agreement (96.94%), followed by ZSP-405b (90.83%) and ZSP-1b (90.52%).
We further performed statistical tests to verify differences between the five models in terms of persuasiveness and dominance. We used a one-way ANOVA and Tukey-HSD posthoc test for each metric, using the statsmodels library [16]. Tables IV and V show the statistical result. First, we observed that model differences affect the persuasiveness ($p<$ 0.001). In detail, pairwise differences are all significant: ZSP-405b versus KD-8b ($p$=0.04), ZSP-8b versus KD-1b ($p$=0.004), and the other eight pairs ($p<$ 0.001) are all significant. Second, we also observed that model differences affect the dominance ($p<$ 0.001). Pairwise differences are all significant except for two pairs: ZSP-405b versus ZSP-8b ($p$=0.246) and ZSP-8b versus KD-8b ($p$=0.891) are statistically insignificant.
TABLE IV: Results of Statistical Test on Persuasiveness
@@ IV-E Results
TABLE V: Results of Statistical Test on Dominance
@@
IV-F Discussion
In open-world games, we believe it is important for merchant NPCs to provide immersive experiences to players. To achive such immersion, developers should consider two essential parts of the merchant: its purpose and position in the game world. We discuss our findings based on these two parts.
@@ IV-F Discussion
Second, the ZSP and KD methods can introduce different merchant personalities. In WoW, there are many merchants in a variety of situations, including differences in race and faction. These factors should be considered when developing merchants to enhance the immersive experience of players. For example, a merchant can use different policies for different players based on their races or factions; a merchant can impose higher prices on rich players or players unfavorable to the merchant. As an early stage, our results imply such distinction is possible using different models. To create an agent favorable to players, developers can opt for models with high agreement rates and low dominance such as ZSP-1b or ZSP-8b. Conversely, if developers are looking for more stubborn merchants, models with low agreement rates and high dominance, like KD-8b, would be a good option.
IV-G Additional case analysis
@@
Hello. I’m looking for Cadet Belt.
@@
The improvisation case indicates that merchants sometimes create spontaneous statements about their stock lists. This error mainly occurred in ZSP-1b, the smallest model without training. An example of this case is shown in the quote following this paragraph. We suspect that this improvisation is a form of “hallucination.” In natural language processing, this term describes instances where language models generate nonsensical or inaccurate text in relation to the provided input sources [17]. Prior studies have also shown that smaller language models are more likely to generate hallucinated outputs [18]. Therefore, developers should be prepared for the possibility that merchants may suggest nonexistent items. This case can be handled through two approaches: mitigation or approval. To mitigate the problem, we recommend that developers implement a retrieval-augmented generation system [19], which encourages language models to produce outputs based on a specific database. For approval, developers should establish a confirmation system to assess the relevance and legitimacy of any newly introduced items by merchants.
[The Conjurer’s Bracers is in WoW Classic, but the Conjurer’s Sigil Clock is not.]
@@
In the arithmetic error case, merchants sometimes struggle with basic calculations involving item quantity or price. This case is typically seen in smaller negotiator such as ZSP-1b, ZSP-8b, KD-1b, and KD-8b. An example of this case is shown in the quote following this paragraph. We attribute such errors to the lack of arithmetic capability in small language models. Previous studies on LLMs also support this speculation because smaller LLMs often struggle with arithmetic tasks [20, 21, 22]. The arithmetic error case may disrupt negotiations and lead to embarrassment for players. Therefore, developers need to address these issues when incorporating smaller LLMs as negotiation modules. One potential solution is to use external calculators to assist with arithmetic calculations.
@@
V Related Work
This section reviews previous studies regarding two passivity issues: communication and pricing. Each section explains existing approaches in games and real-world applications.
V-A Using LLMs to communicate with NPCs
Recently, game developers have started using LLMs in their games to provide more human-like interaction between players and NPCs. Specifically, there is a growing concern about making conversational NPCs with LLMs [23, 24, 25, 26, 27, 28]. For example, Christiansen et al. [25] used LLM-based NPCs to support player interactions within a murder mystery game by helping players interrogate, collect clues, and explore. Gao et al. [27] anthropomorphize chemical entities using LLMs to provide dialogue-based interactions within an educational game. These studies successfully provided human-like communication between NPCs and players.
Despite the successful use of LLMs, previous studies mainly focused on delivering information to the player. In other words, NPCs in those games usually provide their knowledge and reasoning to the player. However, delivering information is not the only form of communication in the real world. As we discussed, this paper addresses another form: negotiation, which frequently appears in transactions between human merchants and customers but is less examined by researchers.
In real-world applications, NLP researchers have investigated the potential of LLMs to negotiate [29, 30, 31]. For example, [29] leveraged LLMs to generate a persuasive dialogue that one may encounter in ordinary life. They made an LLM first choose a persuasion strategy before generating an utterance and achieved state-of-the-art performance. Similarly, [30] suggested a personal negotiation coach based on LLMs and proved their effectiveness in assisting human negotiators. Since these LLM-based studies have not explored NPC design in this context, we attempted to confirm the possibility.
V-B Using LLMs to estimate values
As item prices are usually fixed in a game system, research has not sufficiently focused on predicting item prices based on item descriptions. An active merchant, however, should have the ability to decide item prices in its shop, similar to a real-world situation. Thus, we designed an appraiser module, inspired by studies on estimating prices of real assets.
In real-world applications, researchers have investigated the applicability of LLMs in estimating commodity prices. Early studies attempted to predict prices based on a fixed set of features [32, 33]. For example, [33] used a fixed set of features and a machine-learning method for predicting the price of seasonal goods. However, such models generally do not perform well when predicting prices of out-of-domain goods. To handle unseen goods, researchers utilized latent representations from textual descriptions of items with a language model [34, 35]. For instance, [34] demonstrated that LLMs can predict stock market movements from the earnings report of a company to support information-based investment. Based on these studies, we demonstrated that LLMs can also achieve strong predictive performance in the context of games.
VI Conclusion
@@
We proposed a novel framework called mart, which consists of two LLM-based modules, appraiser and negotiator, which are designed to resolve two issues that make merchant NPCs passive: pricing and communication. We also conducted two experiments comparing different approaches for developing each module to guide game developers in selecting suitable implementations for their deployment conditions. As a result, we demonstrated that our approaches can build an active merchant NPC in WoW Classic. Specifically, we found that one can build strong appraisers and negotiators based on larger LLMs without any additional training. Also, we discovered the possibility of building acceptable appraisers and negotiators with smaller LLMs using supervised finetuning or knowledge distillation methods. Moreover, we found possible concerns that developers can face when building active NPCs. For example, LLMs can show unreliable behavior with in-context learning methods. Also, they may promise giveaways during a negotiation. We expect that these findings can be generalized to other open-world games or other trading NPCs, including auctioneers or pawnbrokers.
@@
Despite these contributions, this study has three limitations. First, our experimental results are based on the Llama LLM family. This may lower the generalizability of our results because different LLMs can perform differently. Second, we simulated player negotiation using GPT-4o. Humans may use different tactics to neutralize the negotiator. Third, we suggested possible solutions for irregular negotiation cases without any tests. Further investigations are required to address these limitations. Nonetheless, we believe that this study can serve as a foundation for further studies on integrating LLMs to create active merchants.
VII Acknowledgements
@@
This work was supported by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT) (No.2021-0-01341, Artificial Intelligence Graduate School Program, Chung-Ang University) and the Chung-Ang University Graduate Research Scholarship in 2024.

### v3

@@ 5.1 Two tested approaches
First, we employed the ZSP method in which an LLM generates negotiation dialogues without being provided any demonstrations. We initially considered a naïve negotiation method without using any specific tactics. This approach relied solely on using the pretrained knowledge of LLMs, which reflects a general understanding of the world; consequently, the generated negotiation strayed significantly from the intended game setting. Thus, we inputted 10 negotiation tactics within the prompt, as shown in Table 2, which are inspired by Cialdini (2001); Orji et al. (2015); Guo and Barnes (2009); Park and Lee (2011). We evaluated three Llama variants as in the appraiser.
@@
Table 2: Tactics used in the input prompt to help in negotiations
@@
Second, we employed knowledge distillation to allow smaller LLMs to achieve comparable performance with lower resource requirements. By transferring the knowledge of a larger model to a smaller model, the distillation reduces computational cost while maintaining performance. We trained student models on a dialogue dataset generated by teacher model. Here, we prompted the teacher model to select an appropriate negotiation tactic before generating each utterance to distill its persuasiveness. To further reduce computational demands during distillation, we applied quantization and low-rank adaptation (LoRA) Hu et al. (2021). Using these methods, we distilled knowledge about negotiation from Llama 3.1 405b to two smaller LLMs: Llama 3.1 8b and 3.2 1b.
5.2 Dataset
@@
We prepared a dataset by generating a negotiation dialogue for each item in the appraiser dataset. To simulate a negotiation dialogue, we used two agents: a merchant and a player. As the merchant, we used Llama 3.1 405B, the teacher model. As the player, we used GPT-4o with input prompts about tactics. We used different models for these two agents to (1) avoid adopting a similar reasoning process and (2) make the player as similar as possible to human beings, as reported in Kwon et al. (2024). To simulate the diversity of real-world customers, we used a temperature value of 1.0 for GPT-4o.
@@
The negotiation procedure was as follows. Before starting a negotiation, both agents received an item description of the negotiation subject as shared information. The two agents were assigned different prices for each item: the player wanted to buy the item at a discount of 10% to 25%. The player started the negotiation with an utterance “Hello. I’m looking for [the item name].” Although greetings and small talk are common in real-world negotiations, we intentionally avoided including such content to ensure that the models learned to focus on goal-directed negotiation behaviors.
@@
After the initial utterance, the two agents engaged in price negotiation until the player decided whether to purchase the item or not. Until the decision was made, the merchant kept persuading the player to buy the item to simulate a real-world merchant. The player could terminate the conversation by mentioning “conversation over.” To avoid long dialogues, we set the maximum number of turns to 15. As a result, we generated 2,943 conversations: 2,616 conversations from the training set items and 327 conversations from the validation set items.
5.3 Evaluation metrics
@@
We compared three ZSP models and two KD models using three evaluation metrics: persuasiveness, dominance, and agreement. First, we measured the persuasiveness of each utterance generated by the tested negotiators. Similar to commercial behavior in the real world, the merchant NPC should effectively persuade players to buy or sell an item at a favorable price. Various tactics are used in an effective negotiation; so, we aimed to evaluate the effectiveness of the negotiator in using 10 different tactics to create persuasive statements. To measure persuasiveness, we used G-Eval Liu et al. (2023), a widely-used evaluation method using LLMs. The method directly asks GPT-4 to evaluate an input text according to given criteria. Specifically, we used a 5-point scale and averaged 20 runs following G-Eval.
@@
Second, we measured the dominance of the merchant over the player during the negotiation. This metric indicates whether the merchant holds more power in the relationship with the player, highlighting the concept of power dynamics Kim et al. (2005). A human merchant usually has the initiative in price negotiation, rather than giving the initiative to the customer. So, as a negotiated price increases, the merchant profits more while the player incurs greater losses. In other words, the negotiation is a type of zero-sum game. Considering such power dynamics, we measured dominance using Equation 4, where $y$, $y^{m}$, and $y^{p}$ indicate the agreed price, retail price, and price desired by the player, respectively. The equation quantifies the ratio between the profit gained by the merchant ($y_{i}-y_{i}^{p}$) and the gap between two agents ($y_{i}^{m}-y_{i}^{p}$).
@@
Third, we defined the agreement rate as the proportion of negotiations in which the merchant and the player reached a mutual settlement on a specific price. While dominance captures how favorable the final price is to the merchant, relying on it alone may lead to misinterpretation. A merchant who persistently offers unreasonably high prices might achieve high dominance in a few successful cases, while failing most negotiations. Therefore, we also consider the agreement rate, which measures how often a mutual settlement is reached across total negotiations.
@@
Table 3: Results of assessing two methods of negotiator. Numbers in parentheses of persuasiveness and dominance indicate the number of utterances and settled negotiations, respectively.
5.4 Results
Table 3 shows the results of our experiments, comparing ZSP and KD methods. The results show that the KD-8b performed the best and successfully used persuasive tactics. The model achieved a score of 3.99 in terms of persuasiveness, followed by ZSP-405b (3.92), ZSP-8b (3.74), KD-1b (3.65), and ZSP-1b (2.95). The ZSP-405b had the strongest dominance (0.47), followed by ZSP-8b (0.42), and KD-8b (0.40). The ZSP-8b had the highest agreement (96.94%), followed by ZSP-405b (90.83%) and ZSP-1b (90.52%).
We further performed statistical tests to verify differences between the five models in terms of persuasiveness and dominance. We conducted a one-way ANOVA to examine whether there were significant differences among groups using the statsmodels library Seabold and Perktold (2010). Also, we conducted a post-hoc analysis using the Tukey-HSD test and $p$-value adjustment to identify group differences. Tables 4 and 5 show the statistical result. First, we observed that model differences affect the persuasiveness ($p<$ 0.001). In detail, pairwise differences are all significant: ZSP-405b versus KD-8b ($p$=0.04), ZSP-8b versus KD-1b ($p$=0.004), and the other eight pairs ($p<$ 0.001) are all significant. Second, we also observed that model differences affect the dominance ($p<$ 0.001). Pairwise differences are all significant except for two pairs: ZSP-405b versus ZSP-8b ($p$=0.246) and ZSP-8b versus KD-8b ($p$=0.891) are statistically insignificant.
@@ 5.4 Results
Table 4: Results of statistical test on persuasiveness
@@
Table 5: Results of statistical test on dominance
5.5 Discussion
In open-world games, we believe that merchant NPCs should contribute to immersive player experiences. To support such immersion, developers need to consider two essential parts of a merchant: its purpose and position within the game world. We discuss our findings with respect to these two aspects.
@@ 5.5 Discussion
Second, the ZSP and KD methods can introduce different merchant personalities. In WoW Classic, there are many merchants in a variety of situations, including differences in race and faction. These factors should be considered when developing merchants to enhance the immersive experience of players. For example, a merchant can use different policies for different players based on their races or factions; a merchant can impose higher prices on rich players or players unfavorable to the merchant. As an early stage, our results imply such distinction is possible using different models. To create an agent favorable to players, developers can opt for models with high agreement rates and low dominance such as ZSP-1b or ZSP-8b. Conversely, if developers are looking for more stubborn merchants, models with low agreement rates and high dominance, like KD-8b, would be a good option.
5.6 Additional case analysis
@@
Hello. I’m looking for a Cadet Belt.
@@
The improvisation case indicates that merchants sometimes create spontaneous statements about their stock lists. This error mainly occurred in ZSP-1b, the smallest model without training. An example of this case is shown in the quote following this paragraph. We suspect that this improvisation is a form of “hallucination.” In natural language processing, this term describes instances where language models generate nonsensical or inaccurate text in relation to the provided input sources Maynez et al. (2020). Prior studies have also shown that smaller language models are more likely to generate hallucinated outputs Carlini et al. (2023). Therefore, developers should be prepared for the possibility that merchants may suggest nonexistent items. This case can be handled through two approaches: mitigation or approval. To mitigate the problem, we recommend that developers implement a retrieval-augmented generation system Lewis et al. (2020), which encourages language models to produce outputs based on a specific database. For approval, developers should establish a confirmation system to assess the relevance and legitimacy of any newly introduced items by merchants.
[The Conjurer’s Bracers is in WoW Classic, but the Conjurer’s Sigil Cloak is not.]
@@
In the arithmetic error case, merchants sometimes struggle with basic calculations involving item quantity or price. This case is typically seen in smaller negotiators such as ZSP-1b, ZSP-8b, KD-1b, and KD-8b. An example of this case is shown in the quote following this paragraph. We attribute such errors to the lack of arithmetic capability in small language models. Previous studies on LLMs also support this speculation because smaller LLMs often struggle with arithmetic tasks Touvron et al. (2023); Grattafiori et al. (2024); OpenAI et al. (2024). The arithmetic error case may disrupt negotiations and lead to embarrassment for players. Therefore, developers need to address these issues when incorporating smaller LLMs as negotiation modules. One potential solution is to use external calculators.
@@
6 Conclusion
@@
We proposed a novel framework named mart, consisting of two LLM-based modules–appraiser and negotiator–designed to resolve two key limitations of passive merchant NPCs: pricing and communication. To support implementation decisions under varying deployment conditions, we conducted experiments comparing multiple implementation approaches for each module. Our results demonstrate that our frameworks can construct an active merchant NPC in WoW Classic using larger LLMs without additional training. We also showed that smaller LLMs can achieve acceptable performance when trained via supervised finetuning or knowledge distillation. Moreover, we identified several concerns developers may face when integrating LLMs into interactive NPCs, including promising unrealisitc giveaways during negotiation. We expect that these findings can generalize to other open-world games and trading NPCs.
@@
Despite these contributions, this study has three limitations. First, our experimental results are based on the Llama LLM family. This may lower the generalizability of our results because different LLMs can perform differently. Second, we simulated player negotiation using GPT-4o. Humans may use different tactics to neutralize the negotiator. Lastly, our discussion centered on misbehavior that may emerge inherently from LLM-based merchants, rather than on adversarial threats like prompt injection by malicious users. Further investigations are required to address these limitations. Nonetheless, we believe this study can serve as a foundation for further studies on integrating LLMs to active merchants.
Acknowledgements
@@
This work was supported by the Institute of Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) [RS-2021-II211341, Artificial Intelligence Graduate School Program (Chung-Ang University)]